OpenStudy (anonymous):

A rectangle with a length of x - 4 and a width of 8 has a perimeter of 34. What is the value of x?

OpenStudy (anonymous):

perimeter = 2(L+b)

OpenStudy (anonymous):

2(x-4 + 8) = 34 x+4 = 17 x= 13